OpenStudy (anonymous):

Which of the following is an example of the inverse of an exponential function

OpenStudy (anonymous):

y= x/4 y=4x y=2^x y=log2x

OpenStudy (aravindg):

y=log 2x

OpenStudy (aravindg):

the inverse of exponential function is logarithmic function

OpenStudy (anonymous):

I thought so just wasn't sure..

OpenStudy (aravindg):

:) now u r sure when i am here! :)

OpenStudy (anonymous):

Yupp lol :D Thank you!